### v1

# Semantic Segmentation Algorithm Based on Light Field and LiDAR Fusion

Abstract
Semantic segmentation serves as a cornerstone of scene understanding in autonomous driving but continues to face significant challenges under complex conditions such as occlusion. Light field and LiDAR modalities provide complementary visual and spatial cues that are beneficial for robust perception; however, their effective integration is hindered by limited viewpoint diversity and inherent modality discrepancies. To address these challenges, the first multimodal semantic segmentation dataset integrating light field data and point cloud data is proposed. Based on this dataset, we proposed a multi-modal light field point-cloud fusion segmentation network(Mlpfseg), incorporating feature completion and depth perception to segment both camera images and LiDAR point clouds simultaneously. The feature completion module addresses the density mismatch between point clouds and image pixels by performing differential reconstruction of point-cloud feature maps, enhancing the fusion of these modalities. The depth perception module improves the segmentation of occluded objects by reinforcing attention scores for better occlusion awareness. Our method outperforms image-only segmentation by 1.71 Mean Intersection over Union(mIoU) and point cloud-only segmentation by 2.38 mIoU, demonstrating its effectiveness.
I Introduction
As a fundamental task in computer vision, semantic segmentation is crucial for a wide range of applications, including autonomous driving [1], road detection [2], and medical image processing [3]. Existing semantic segmentation methods can be divided into image-based semantic segmentation[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and LiDAR-point-cloud-based semantic segmentation[18, 19, 20, 21, 22, 23, 24, 25] according to different types of input data. Image-based semantic segmentation aims to assign a specific semantic label to each pixel in the image, while LiDAR-point-cloud-based semantic segmentation aims to assign a specific semantic label to each point in the point cloud.
Images provide rich color and texture information, but lack accurate 3D spatial structure and are highly sensitive to lighting conditions. In contrast, LiDAR point clouds offer precise spatial geometric data and are less affected by lighting due to their reliance on wavelengths but lack color and texture information. Therefore, effectively fusing images and point clouds is crucial to leverage the strengths of both modalities, enabling more accurate and reliable semantic segmentation.
Recent works, such as 2DPASS [26] and Mseg3D [27], have fused LiDAR point clouds with images to improve 3D segmentation accuracy. CMNeXt [28] integrates LiDAR point cloud benefits for 2D image segmentation. These fusion methods have advanced both 2D and 3D semantic segmentation. However, despite using multiple modalities, they ultimately segment based on a single modality, failing to fully exploit both. When LiDAR point clouds are fused with images but only the images are segmented, the disparity in density between them reduces fusion effectiveness. Sparse point clouds often hinder image segmentation, rather than helping. Moreover, although current multimodal datasets include both cameras and LiDAR, the sparsity of LiDAR point clouds provides limited supplementary information for occluded objects. Additionally, the multiple cameras in existing multimodal datasets often lack overlapping fields of view, making it difficult to effectively complement occluded regions.
Light field images, with multiple viewpoints and significant overlap, provide enhanced occlusion perception by capturing more angular information. By recording different visible parts of occluded objects from various perspectives, which provide more comprehensive cues for occluded content. For example, UrbanLF [4] provides multiple sub-aperture images for richer angular data. However, existing datasets captured using camera arrays offer a large baseline, which is beneficial for exploiting angular information in outdoor scenes, they typically lack semantic annotations, making them unsuitable for segmentation tasks. In contrast, datasets collected using light field cameras provide annotated data, yet their small baseline limits angular diversity and reduces the effectiveness of multi-view fusion. Moreover, most of these datasets only annotate the central view, further restricting methods that aim to leverage full-view light field information.
To address the issues identified above, we constructed TrafficScene [29], the first dataset with semantic annotations that includes both light field images and LiDAR point cloud data. Unlike previous datasets, all viewpoints of the light field are annotated, enabling effective information supplementation for occluded and small objects through multi-view consistency. To integrate light field data and point cloud data effectively, we propose a novel light filed and point cloud fusion segmentation algorithm that aims to fully leverage the complementary strengths of both modalities through simultaneous segmentation of the light filed and point cloud data. To overcome the challenge of poor light filed image segmentation caused by the varying sparsity between point cloud and light filed image during fusion, a pixel-point feature fusion interpolation module is proposed. This module interpolates the features of point clouds projected onto the light filed image plane and subsequently fuses them, thus mitigating the negative impact of sparse point clouds on light filed image segmentation. To enhance the recognition of occluded objects, we introduce a depth difference perception module, which leverages depth information to perceive occlusions.
The major contributions are as follows:
1. TrafficScene, the first multimodal dataset for semantic segmentation that incorporates light field modalities. Captured using a unique 3×3 camera array with a 30 cm baseline, TrafficScene provides comprehensive semantic annotations across all light field viewpoints, enabling effective multi-view information utilization.
2. We propose the first simultaneous light filed and point cloud segmentation method, Multimodal Light Filed Point Cloud Fusion Segmentation Method (Mlpfseg), it enhances the full integration of point clouds and images and improves the perception of occluded objects through the pixel-point feature fusion interpolation module and the depth difference perception module. Compared to single-image semantic segmentation, the mIoU improves by 1.71. Compared to light field image segmentation, the mIoU improves by 2.37, and compared to multimodal 3D semantic segmentation, the mIoU improves by 2.38.
The rest of this paper is organized as follows. Section II summarizes related works on semantic segmentation datasets and semantic segmentation methods. The proposed approach is detailed in Section III. Experiments, including comparisons, ablation studies and visualization are given in Section IV. Finally, we conclude the paper in Section V.
II Related Works
In this section, we introduce existing light field semantic segmentation datasets and semantic segmentation methods. In this section, we introduce existing semantic segmentation datasets, image semantic segmentation methods, light field semantic segmentation methods, point cloud semantic segmentation methods and multimodal fusion semantic segmentation methods.
II-A Semantic Segmentation Dataset
Among existing semantic segmentation datasets, image-based datasets[30, 31, 32] typically rely on a single perspective, which limits their ability to capture occluded or small objects effectively. Light field image semantic segmentation datasets[4] leverage multiple viewpoints to partially address occlusions; however, they are usually collected using light field cameras with narrow baselines between views, and only the central view is annotated with semantic labels. This limits their effectiveness in supplementing occluded and small object information. Point cloud semantic segmentation datasets[33, 34] provide accurate 3D spatial information and are effective in recognizing occluded objects, but they are inherently sparse, making it difficult to achieve high segmentation accuracy. Existing multimodal semantic segmentation datasets[35, 36, 37, 38, 39] utilize cross-modal fusion to improve recognition of small and occluded objects to some extent. However, since they also rely on a single viewpoint, their ability to comprehensively supplement occluded or small object information remains limited.
II-B Image Semantic Segmentation Methods
Early semantic segmentation algorithms relied on manual feature extraction [6], but with deep learning advancements, recent methods use deep learning for feature extraction. The Fully Convolutional Network (FCN)[7] pioneered deep learning in image segmentation with an end-to-end network using convolutional downsampling and bilinear interpolation. The Pyramid Scene Parsing Network (PSPNet) [8] introduced pyramid pooling for varied object sizes, enhancing precision across scales. DeepLabV3 [9] improved multi-scale object segmentation with various sampling rates for multi-scale contextual information. Recent transformer-based approaches have furthered accuracy in image semantic segmentation. OCRNet [10] employed attention mechanisms for enhanced precision, and Mask2former [11] used masked attention for precise segmentation of smaller objects. SegFormer [12] integrated convolution with MLP for improved object connectivity. Despite theseadvances, single-image semantic segmentation struggles in occlusion or color similarity scenarios due to its limited spatial geometric information capture.
II-C Light Field Semantic Segmentation Methods
Recent algorithms in light field semantic segmentation utilized spatial and angular data to enhance precision, especially in occlusion detail. Chen et al. [5] used CNNs with an angular model and ASPP in light field image segmentation. Sheng et al. [4] stacked images to utilize information from different viewpoints, while Cong et al. [13] applied attention mechanisms and depth maps for enhanced feature extraction. Zhang et al. [14] integrated various perspectives using feature rectification and fusion modules. [15] proposes LF-IENet++, a light field semantic segmentation network that effectively handles multi-baseline disparities via feature integration and propagation. Despite these advancements, the effectiveness of these methods is constrained by the small baseline of light field cameras, limiting the additional information from different viewpoints.
II-D Point Cloud Semantic Segmentation Methods
Unlike images, semantic segmentation of point clouds aims to assign specific categories to each point cloud. Point clouds have gained various uses in autonomous driving due to their unique and accurate three-dimensional spatial structure. There are three main approaches to semantic segmentation of point clouds: directly operating on points, segment after projection, and based on voxels. Methods that directly operate on points, such as PointNet [18] and PointNet++ [19], utilize multi-layer perceptrons to extract features from the point clouds, combining local and global features to classify each point. JSNet++[25] boosts 3D point cloud segmentation via dynamic convolutions and spatial-channel correlation modeling. Projection-based methods map point cloud data onto 2D image coordinates via projection, then apply classic 2D convolutional network architectures for segmentation. Typical projection methods include spherical projection [20] and bird’s-eye projection [21]. A notable algorithm in this category is SqueezeSeg [22], which first converts point clouds into front views using spherical projection before performing segmentation. Recently, voxel-based methods have gained attention. These methods divide the 3D world into voxels and use 3D convolution operations within each voxel block, ultimately obtaining segmentation results for each point in each voxel through upsampling. MinkowskiNet [23] introduced sparse convolution, which efficiently processes high-dimensional sparse data, while SPVCNN [24] divides voxels into regular blocks and applies sparse convolution operations. Methods focused solely on point clouds for 3D semantic segmentation face challenges due to their inherent sparsity. This sparsity leads to a lack of information about small and occluded objects, resulting in a lower segmentation accuracy for these objects.
II-E Multimodal Fusion Semantic Segmentation Methods
Recent algorithms combine images and point clouds to boost segmentation accuracy, generally falling into two categories: those that fuse both modalities but segment only the image, and those that fuse both and segment only the point cloud. In image segmentation, methods like CMNeXt [28] project the point cloud onto the image plane for feature fusion, but the sparsity of point cloud data limits effectiveness. For point cloud segmentation, there are two fusion approaches: data-level and feature-level. In data-level fusion, methods like FuseSeg [40] project the RGB image onto the point cloud’s spherical projection and segment the point cloud, though this can lose intrinsic data structure. Feature-level fusion, such as PMF [41] and 2DPASS [26], extracts multi-scale features from both modalities and fuses them. MSeg3D [27] advances fusion by introducing cross-modal attention, yielding the best results in multimodal fusion methods.
Although recent multimodal fusion semantic segmentation approaches leverage multiple modalities for feature fusion, they typically produce segmentation results for a single modality. This limitation hinders the effective integration of complementary information from both images and point clouds, thereby constraining segmentation performance, particularly in the presence of occluded and small-scale objects. Although multimodal fusion outperforms single-modality point cloud approaches in these scenarios, the improvement remains limited due to the inherent perspective constraint of single-camera imaging, which provides only partial scene information. Therefore, the fusion of light field images and LiDAR point clouds can further enhance the segmentation accuracy of occluded and small objects. To address the challenges of insufficient fusion caused by the difference in spatial density between light field images and point clouds, as well as the limited capability of existing networks to accurately segment occluded and small objects, we propose a novel methods: Mlpfseg. Specifically, we introduce the Point-Pixel Feature Fusion Module (PFFM) to effectively integrate sparse point cloud data with dense light field image. In addition, the Depth Difference Perception Module (DDPM) is designed to enhance the recognition of occluded objects by leveraging depth inconsistency cues.
(a) The light field image of the camera
(b) Annotations of the light field image
(c) The point cloud data
(d) The RGB projection of the image
(e) Annotation of the point cloud
Figure 1: Examples of the data we collected.
(a) Collectionvehicle
(b) Collectionequipment
Figure 2: Multimodal acquisition system.
III Proposed Method
In this paper, we introduce the first multimodal dataset incorporating both light field images and LiDAR point clouds. The light field images are captured using a camera array with a large baseline, providing multiple perspectives that complement each other for better occlusion handling. Annotations are provided for all viewpoints to facilitate full integration. Based on this dataset, we propose a fusion algorithm: Multimodal Light Filed Point Cloud Fusion Segmentation Method (Mlpfseg). Through Point-Pixel Feature Fusion Module(PFFM) and Depth Difference Perception Module(DDPM), we enhance the fusion process, significantly improving segmentation accuracy, especially for occluded objects. These advancements will be discussed in the following sections.
III-A Multimodal TrafficScene Dataset
Figure 3: The proportion of annotated pixels (y-axis) per class (x-axis) in TrafficScene, Cityscapes [30], UrbanLF [4].
TrafficScene represents the first multimodal dataset that offers jointly annotated point clouds and full-viewpoint light field images. The following will introduce the collection, calibration, annotation, and statistical analysis of the multimodal dataset. A multimodal dataset of 5607 light field images and 623 frame point clouds was constructed in traffic scenarios using a 3×3 FLIR BFS-PGE 16S2C camera array[42] with a 30 cm baseline and CH128X1 LiDAR[43]. Each frame of the point cloud has more than 60,000 points. The multimodal data acquisition system is shown in Fig. 2. This setup enhances angular diversity and depth perception, aiding in the detection of small or occluded objects and improving semantic segmentation accuracy in complex outdoor scenes. Compared to conventional light field systems, it reduces rain artifacts and improves edge detection. Accurate spatial alignment between modalities is ensured via pairwise calibration. Our aerial perspective captures varied scenes during peak traffic hours, yielding high-resolution (1440×1080) multimodal data. Additional details regarding the dataset are provided in the supplementary material. The dataset is available for download at: https://pan.baidu.com/s/1pY9uY0JP52IeJXmgvfeVcQ.
All light field viewpoints are distortion-corrected and annotated with 15 semantic categories using CVAT[44] As shown in Fig. 3. Annotations are projected onto the LiDAR data and manually refined, making this the first multimodal dataset with aligned semantic labels across both modalities. Our dataset encompasses five representative traffic scenarios—parking lots, urban roads, vegetated roads, roadside areas adjacent to buildings, and roads with obstacles—as well as five types of common traffic participants, including cyclists, pedestrians, buses, cars, and bicycles. Background classes like vegetation and roads dominate, while rich annotations of traffic-specific elements enhance urban scene understanding, offering significant potential for autonomous driving and intelligent transportation systems.
Figure 4: Internal structure of multimodal light field point cloud fusion segmentation network. It mainly consists of two parts: point-pixel interpolation fusion module (PFFM) and depth difference perception module (DDPM).
III-B Semantic Segmentation Algorithm Based on Multimodal Data Fusion
Fig. 4 provides the diagram of the proposed Multimodal Light Field Point Cloud Fusion Segmentation Method (Mlpfseg). Mlpfseg consists of two branches : the light field image branch and the point cloud branch, which are specifically designed for the extraction of image features and point cloud features, respectively.
For the light field image branch, the input consists of light field images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, where each image has the size of $\mathbb{R}^{3\times H\times W}$ and $n$ denotes the number of input camera viewpoints. To extract the image features, we employ a weight-shared HRNet-48[45], which enables efficient multi-scale feature representation and enhances the model’s ability to capture detailed spatial information, obtaining the corresponding viewpoint features $\{F_{img1},F_{img2},\cdots,F_{imgn}\}$. The size of each viewpoint feature is $\mathbb{R}^{c_{img}\times h\times w}$, here $c$ represents the number of feature channels; $h$ denotes the height of the feature map; $w$ represents the width of the feature map.
For the point cloud branch, the input point cloud is denoted as $P_{point}\in\mathbb{R}^{N\times 4}$, where $N$ represents the number of points in the point cloud, and 4 corresponds to $\{x_{i},y_{i},z_{i},r_{i}\}$, which represent the 3D spatial coordinates and the refractive index of the $i$-th point, respectively. The point cloud is divided into voxels, and the coordinates of the $i$-th point assigned to the $k$-th voxel are given by $Voxel_{k}=\left\{\left(\left\lfloor\frac{x_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{y_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{z_{i}}{r_{l}}\right\rfloor\right)\right\}\in\mathbb{R}^{N\times 3}$. Here, sparse convolution is utilized to extract voxel features through SPVCNN[24], resulting in $F_{voxel}^{l}\in R^{N_{1}\times c_{p}}$, where $F_{voxel}^{l}$ represents the features of the voxels extracted at the $l$-th layer, $N_{1}$ denotes the number of non-empty voxels, and $c_{p}$ represents the number of feature channels per voxel. After obtaining the voxel features at the $l$-th layer, the corresponding features for each point can be derived through interpolation:
$F_{point}^{l}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{voxel}^{l}\quad,$ (1)
where $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$ and $w_{i}=\frac{1}{d(p,v_{i})+\epsilon}$. Here, $p$ represents the point to be interpolated; $v_{i}$ denotes the index of the nearest neighboring voxel; $d(p,v_{i})$ represents the distance between the point and the center of the voxel. The normalized weight is denoted by $\hat{w}_{i}$. The interpolated point features are represented as $F_{point}^{l}\in\mathbb{R}^{N\times c_{p}}$, where $N$ is the number of points; $c_{p}$ is the number of feature channels per point. A small constant $\epsilon=1\times 10^{-8}$ is added to prevent division by zero.
After the extraction of the features, the Point-Pixel Feature Fusion Module (PFFM) is proposed to fuse the image features $\{F_{img1},F_{img2},\cdots,F_{imgn}\}$ and the features of the voxels $F_{voxel}^{i}$, which will be discussed in detail in the following subsection Point-Pixel Feature Fusion Module (PFFM). In PFFM, the sparse characteristics of point clouds will be completed, and the fused feature map $\hat{F}_{fused}$ is obtained.
On this basis, a Depth Difference Perception Module (DDPM) is proposed, with the input predicted depth map for each image $D_{pred}$ and the sparse depth map $D_{sparse}$ presenting the depth values for 3D LiDAR coordinates projected onto the image plane. By utilizing depth difference perception, we obtain the attention score map $\hat{D}_{diff}$ for the occluded objects and send it into the PFFM to optimize the representation in the sparse point cloud module. The detailed description of DDPM will be presented in subsection Depth Difference Perception Module (DDPM).
Ultimately, by inputting the $F_{voxel}^{l}$ for each layer , Mseg3D [27], including multi-scale feature extraction modules and context information fusion modules, and the segmentation head are applied to obtain the fused output $\hat{y}_{img}$ in the image branch and the output $\hat{y}_{point}$ in the point cloud branch.
III-B1 Point-Pixel Feature Fusion Module (PFFM)
After obtaining the point-level features $F_{point}^{i}$ for the $i$-th point cloud in point cloud branch, we project them onto the image plane. Given the original coordinates $\{x_{i},y_{i},z_{i}\}$ of the $i$-th point cloud, the projected coordinates on the image plane are computed as:
$\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}^{\top}=\frac{1}{z_{i}}\times\mathbf{K}\times\mathbf{T}\times\begin{bmatrix}x_{i}\\ y_{i}\\ z_{i}\\ 1\end{bmatrix}^{\top}$, (2)
where $\mathbf{K}\in\mathbb{R}^{3\times 4}$ is the camera intrinsic matrix; $\mathbf{T}\in\mathbb{R}^{4\times 4}$ is the camera extrinsic matrix. Here, $u_{i}$ and $v_{i}$ are the coordinates of the projected point on the image plane obtained through perspective projection. Since the feature map size is smaller than the original image size due to feature extraction by HRNet-48, the corresponding coordinates on the feature map are given by $u_{i}^{\prime}=u_{i}\times\frac{h}{H}$ and $v_{i}^{\prime}=v_{i}\times\frac{w}{W}$. The projected features on the image plane are denoted as $F_{point}^{img}$.
Considering the characteristics of point cloud projections on the image, where the point cloud is densely concentrated in the central region of the image plane, and there are no point clouds in the upper and lower areas, we design the following interpolation method. First, we find the minimum bounding rectangle $M$ for all the projected point clouds on the image plane and let the feature points that can be projected within this rectangle be denoted as $P_{in}^{point}\in\mathbb{R}^{N\times c_{p}}$. There are no matching point clouds for pixels outside the rectangle. At this point, the point cloud feature map projected onto the image plane is given by $F_{point}^{img}=\left[P_{in}^{point},0\right]$. For the pixel positions within the rectangle that do not have corresponding point cloud projections, we perform feature interpolation as follows.
Let the pixel coordinates of the valid projected points $P_{in}^{point}$ be: $(x_{min},y_{min})=\min(x_{indices}),\min(y_{indices})$ and $(x_{max},y_{max})=\max(x_{indices}),\max(y_{indices})$. We define the range of the minimum bounding rectangle $M$ as $x_{min},x_{max}]\times[y_{min},y_{max}]$.
A grid of coordinates $(x,y)$ is generated within the rectangle $M$, where the points with assigned values (i.e., those that have point cloud masks) are labeled as:
$\text{mask}(x,y)=\begin{cases}1&(x,y)\in\{(x_{indices},y_{indices})\}\\ 0&(x,y)\notin\{(x_{indices},y_{indices})\}\end{cases}^{\top},$ (3)
The set of assigned points is: $\{(x,y)\mid\text{mask}(x,y)=0\}$.
For each unassigned point $(x,y)$, we find its three nearest valid points $\{(x_{i},y_{i})\}_{i=1}^{3}$, calculate the interpolation weights, where the weights are inversely proportional to the distance between the unassigned point and The valid points are: $w_{i}=\frac{1}{d_{i}+\epsilon}$, where $d_{i}=\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}}$. We normalize the weights as $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$. The interpolated features are then:
$\hat{F}_{point}^{img}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{point}^{img}(x_{i},y_{i})\quad,$ (4)
Thus, the final interpolated feature for $(x,y)$ is:
$\hat{F}^{img}_{point}(x,y)=\begin{cases}{\scriptstyle\sum_{i=1}^{3}\frac{d_{i}+\epsilon}{\sum_{j=1}^{3}d_{j}+\epsilon}\cdot F^{img}_{point}(x_{i},y_{i}),\quad\text{mask}(x,y)=1}\\[5.0pt] F^{img}_{point}(x,y),\quad\text{mask}(x,y)=0\end{cases}^{\top},$ (5)
This gives us the point cloud projection in the feature map of the image plane within the bounding rectangle $M$. For regions outside the rectangle, we fill them with the corresponding image feature $F_{img}$ to obtain the complete point cloud projection feature map filled point $F_{fill\_point}$ is given by $F_{img}(x,y)$ if $(x,y)\notin M$, and $\hat{F}_{point}^{img}(x,y)$ if $(x,y)\in M$. Since the point cloud and image data features are in different spaces and have different network structures, they are not in the same feature space. To facilitate alignment of the point cloud projection feature map with the image feature space for full fusion, we design an alignment loss function(Pixel2point Loss). This loss function minimizes the difference between the feature spaces of the point cloud and image, enabling better fusion of the two. The alignment loss uses Mean Squared Error (MSE) to measure the difference between the point cloud feature map and the image feature map:
$\mathcal{L}_{align}=\frac{1}{N}\sum_{(x,y)}\left\|F_{fill\_point}(x,y)-F_{img}(x,y)\right\|_{2}^{2}\quad,$ (6)
After obtaining the complete point cloud projection feature map $F_{fill\_point}$, we concatenate it with the image feature map $F_{img}$ to create the initial fused feature map:
$F_{fused}=concat\left([F_{fill\_point},F_{\text{img}}]\right)\quad,$ (7)
Here, $concat(\cdot)$ denotes the concatenation operation.
We then apply a self-attention mechanism to the initial fused feature map to further refine the fusion, resulting in the final fused feature map $\hat{F}_{fused}$. The goal of the self-attention mechanism is fourfold: 1. To enhance the attention of the image feature map to itself. 2. To enhance the attention of the point cloud projection feature map to itself. 3. To allow the image feature map to gather useful information from the point cloud projection feature map. 4. To allow the point cloud projection feature map to gather useful information from the image feature map. We generate the Query, Key, and Value matrices from the fused feature map $\hat{F}_{fused}$: $Q=F_{fused}\cdot W_{Q}\in\mathbb{R}^{C_{q}\times(h\times w)}$, $K=F_{fused}\cdot W_{K}\in\mathbb{R}^{C_{k}\times(h\times w)}$, and $V=F_{fused}\cdot W_{V}\in\mathbb{R}^{C_{v}\times(h\times w)}$, where $W_{Q}\in\mathbb{R}^{C\times C_{q}}$, $W_{K}\in\mathbb{R}^{C\times C_{k}}$, and $W_{V}\in\mathbb{R}^{C\times C_{v}}$ are learnable projection matrices. The attention weights are calculated based on the similarity between the Query and Key:
$Attention(Q,K,V)=Softmax\left(\frac{Q^{T}K}{\sqrt{C_{k}}}\right)V^{T}\in\mathbb{R}^{(H\times W)\times C_{v}},$ (8)
The attention weights are then applied to the Value to get the final fused feature map $\hat{F}_{fused}$:
$\hat{F}_{fused}=\text{LayerNorm}\left(\text{Reshape}\left(Attention(Q,K,V)\right)\right),$ (9)
Finally, the output of the fused feature map $\hat{y}_{fused}$ is obtained by upsampling $\hat{F}_{fused}$, and the loss is computed by comparing it with the ground truth:
$L_{fused\_img}=CE(\hat{y}_{fused},y_{gt})\quad,$ (10)
III-B2 Depth Difference Perception Module (DDPM)
For occluded objects, fusing point cloud and image data may lead to conflicting features, as a single perspective alone cannot effectively supplement the information of the occluded objects. Thus, it is essential to assist the network in identifying these ’conflicting’ features. We hypothesize that when an object is occluded in the image but not in the point cloud, the depth of the object in the image should differ from its depth in the point cloud. Similarly, if an object is visible in the image but occluded in the point cloud, the depth of the object in the image should also differ from that in the point cloud. Therefore, we propose leveraging depth priors to guide the network in detecting regions with depth discrepancies.
For the input light filed images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, we use Zoe[46], a state-of-the-art deep learning model that leverages convolutional neural networks to predict depth maps from monocular images by learning spatial features and contextual information, to estimate the absolute depth of the images. Through Zoe, we obtain the predicted depth map for each image, denoted as $D_{pred}$. Meanwhile, the LiDAR can acquire 3D spatial coordinates for each point in the point cloud, as described in 2. The 3D coordinates of each LiDAR point can be projected onto the image plane to obtain the corresponding pixel coordinates $(u_{i},v_{i})$ and depth value $z_{i}$. These depth values are then filled into the corresponding positions on the image plane, generating the sparse depth map $D_{sparse}$. For each projected At this point, we have $D_{sparse}(u_{i},v_{i})=z_{i}$ for locations with projection points, and $D_{sparse}(u_{i},v_{i})=0$ for locations without projection points.
We only consider the positions with projection points. The difference map $D_{\text{diff}}$ is obtained by comparing the real sparse depth map and the predicted depth map, serving as the criterion for identifying occluded regions.
Considering that the accuracy of depth prediction from images dramatically decreases as the actual depth increases, we apply logarithmic smoothing to exaggerate the difference at close distances while minimizing the difference at farther distances. That is:
$D_{diff}(i,j)$ $=\log(D_{pred}(i,j)+\epsilon_{i,j})-\log(D_{gt}(i,j)+\epsilon_{i,j})$
$\quad\text{for}\quad(i,j)\in\mathcal{V}_{0},\quad\text{for}\quad(i,j)\notin\mathcal{V}$ (11)
where $\mathcal{V}$ represents the region with LiDAR projection points and $\epsilon=1\times 10^{-8}$. To map the depth differences to the same feature space as the network, we apply a two-layer convolutional neural network to $D_{diff}$:
$\hat{D}_{diff}=Conv(Conv(D_{diff}))\quad,$ (12)
The resulting difference feature map $\hat{D}_{diff}$ is then added to the attention map from Point-Pixel Feature Fusion Module, with the goal of guiding the network to focus on the occluded areas as much as possible, resulting in the final attention map:
$\text{Attention}_{final}=\hat{D}_{diff}+Attention(Q,K,V)\quad,$ (13)
This is then incorporated into the network to enhance the perception of occluded objects.
For the image branch, we use the fused feature loss $L_{fused\_img}$ and the image features $F_{img}$, which are upsampled to obtain the fused output $\hat{y}_{img}$. The loss is calculated as:
$L_{img}=CE(\hat{y}_{img},y_{gt})\quad,$ (14)
For a single image input, the total loss is:
$L_{img\_total}=L_{img}+L_{img\_lovasz}+L_{fused\_img}+\mathcal{L}_{align},$ (15)
For multiple image inputs, the total loss becomes:
$L_{img\_total}=L_{img\_center}+L_{img\_lvcenter}+L_{img\_lovasz}+\mathcal{L}_{align}\\ +L_{fused\_img}+\alpha_{1}\cdot\sum_{i=1}^{n}L_{img\_i}+\alpha_{2}\cdot\sum_{i=1}^{n}L_{img\_lvi},$ (16)
For the point cloud branch, we combine point-level features with image features and apply the segmentation head for point cloud predictions. The voxel features $F_{\text{voxel}}^{l}$ are also upsampled for segmentation. The point cloud and voxel losses are calculated as:
$L_{point}=CE(\hat{y}_{point},y_{point})+L_{lovasz}(\hat{y}_{point},y_{point}),$ (17)
$L_{voxel}=CE(\hat{y}_{voxel},y_{voxel})+L_{lovasz}(\hat{y}_{voxel},y_{voxel}),$ (18)
The total loss for the point cloud is:
$L_{point\_total}=L_{point}+L_{voxel}\quad,$ (19)
The overall network loss is:
$L_{total}=L_{img\_total}+L_{point\_total}\quad,$ (20)
IV Experimental Results
To validate the effectiveness of our proposed the multimodal fusion-based segmentation approach, we conduct experiments on the TrafficScene dataset. All experiments were conducted on a server with an Intel 6330 CPU, 1.0 TB memory, Ubuntu 22.04.5 and CUDA version 12.2. The dataset is split into training, validation, and test sets at a 7:1:2 ratio (3924/594/1116 light filed images, 436/66/124 point clouds), using stratified sampling to balance category distributions.
IV-A Semantic segmentation algorithm based on multimodal data fusion
IV-A1 Experimental Results
To validate the effectiveness of our dataset and multimodal fusion method, we conducted extensive experiments. First, we assessed the dataset’s validity by applying established image and light-field semantic segmentation methods, evaluating performance using mean intersection over union (mIoU).
TABLE I: Quantitative results for image and point cloud semantic segmentation on TrafficScene. Values in parentheses show improvements over previous methods. Red font indicates the state-of-the-art, while blue represents the second-best result.
Method Image Point Cloud Light Field Image mIoU Point Cloud mIoU
FCN [7] $\checkmark$ $\times$ $\times$ 81.23
PSPNet [8] $\checkmark$ $\times$ $\times$ 81.27
DeepLabV3 [9] $\checkmark$ $\times$ $\times$ 80.05
OCRNet [10] $\checkmark$ $\times$ $\times$ 82.27
Mask2Former [11] $\checkmark$ $\times$ $\times$ 82.18
SegFormer [12] $\checkmark$ $\times$ $\times$ 83.26
PSPNet_LGA[29] $\times$ $\times$ $\checkmark$ 81.67
CMNeXt [28] $\times$ $\times$ $\checkmark$ 83.26
MinkowskiNet [23] $\times$ $\checkmark$ $\times$ 84.36
SPVCNN [24] $\times$ $\checkmark$ $\times$ 85.67
2DPASS [26] $\checkmark$ $\checkmark$ $\times$ 70.89
PMF [41] $\checkmark$ $\checkmark$ $\times$ 74.96
Mseg3D [27] $\checkmark$ $\checkmark$ $\times$ 90.00
Baseline $\checkmark$ $\checkmark$ $\times$ 81.32 90.00
Mlpfseg (one view) $\checkmark$ $\checkmark$ $\times$ 85.23(+3.91) 91.50(+1.50)
Mlpfseg (light filed images) $\times$ $\checkmark$ $\checkmark$ 84.97(+3.75) 92.38(+2.38)
Figure 5: mIoU for PSPNet, PSPNet LGA and Mlpfseg on small objects across all viewing angle
(a) Mlpfseg can correctly identify occluded bike
(b) Mlpfseg can correctly identify occluded pedestrian
Figure 6: Visualization of the results of different algorithms for occluded objects
Figure 7: Qualitative results of Mlpfseg on the test set of TrafficScene. Our baseline has a higher error recognizing small objects and partially occluded objects.
For point cloud segmentation, we benchmarked two pure point cloud methods, SPVCNN[24] and MinkowskiNet[23], alongside three multimodal fusion-based methods: PMF[41], 2DPASS[26], and Mseg3D[27]. Additionally, we evaluated our proposed Mipfseg and Mlpfseg. The final results are presented in TABLE I.
The results show that attention-based methods outperform convolution-based ones in single-image segmentation, as attention mechanisms adaptively focus on crucial regions. CMNeXt[28] achieves the highest single-image mIoU of 83.15, while incorporating additional sub-aperture images in light-field-based methods further improves segmentation, surpassing the state-of-the-art (SOTA) with a mIoU of 83.61. This highlights the contribution of multiple viewpoints to central-perspective segmentation quality. For point cloud segmentation, Mseg3D[27] outperforms other methods, demonstrating the benefits of image-assisted segmentation. However, projection-based methods like PMF[41] suffer from 3D structure loss, resulting in inferior performance compared to pure point cloud methods such as SPVCNN[24] and MinkowskiNet[23].
Our proposed Mlpfseg surpass existing SOTA methods in both image and point cloud segmentation. By interpolating image features, they alleviate the sparsity issues caused by point cloud projection. The attention mechanism enhances feature extraction for the image branch, while the depth-difference perception module improves occluded object segmentation, leading to overall performance gains in both modalities.
For both image and point cloud branches, better feature extraction from the image branch also facilitates the fusion of point cloud features. In contrast, the improved feature extraction from the point cloud branch enhances the fusion for the image branch. This demonstrates that with effective information complementarity between image and point cloud data, both modalities can achieve better perception results.
As illustrated in Fig.5, our approach significantly improves mIoU for small objects like bicyclists, pedestrians and traffic cones. The combination of multiple perspectives and multiple modalities provides more complete information on small objects and improves the segmentation accuracy of small objects.
As illustrated in Fig.6, compared with other light field semantic segmentation methods, our method performs better in the cases of missed segmentation and incorrect segmentation of occluded objects. This is because we fully consider the occlusion problem caused by the occlusion relationship of different modalities and use the ddpm module to perceive the occluded objects.
The point cloud visualization of Mlpfseg on the test set is shown in Fig. 7. The red regions highlight the mispredicted points. The visualization reveals that when the bus is partially occluded, our Mlpfseg method significantly outperforms the baseline. Additionally, for small objects such as pedestrians, our method demonstrates improved accuracy. This performance gain can be attributed to our comprehensive fusion of image and point cloud data, along with the design of the Depth Difference Perception Module (DDPM), which enhances the network’s ability to recognize occluded objects.
IV-A2 Implementation details
Image Semantic Network Implementation: For the input images (1080×1440), we apply several augmentation techniques, including random horizontal flipping (50%), color jittering, JPEG compression noise (quality range [30, 70]), and random cropping (60%-75%). The image segmentation network uses HRNet-W48 as the backbone to extract multi-scale features. We fine-tune a pre-trained HRNet-W48 model on ImageNet, freezing the first three stages. After feature extraction across four stages, the image features $F_{\text{img}}\in\mathbb{R}^{c_{\text{img}}\times h\times w}$ (where $c_{\text{img}}=48$) are fused and passed through a modified FCN head for segmentation. For Mipfseg, we use 3 cameras. Point Cloud Semantic Network Implementation: The point cloud network uses a modified UNet3D architecture with an improved voxel feature extractor. The voxelization occurs within the Cartesian space $x,y,z\in[-50,6,-7]$ to [50, 106, 11], with a resolution of 0.1m and a max of 5 points per voxel. The UNet3D structure applies 8× downsampling and upsampling with a channel scaling factor of 2. After encoding, the point cloud features are $F_{\text{voxel}}^{l}\in\mathbb{R}^{N_{1}\times c_{p}}$, where $c_{p}=48$. Other configurations follow Mseg3D [27]. Training Configuration: Both networks are trained end-to-end using the Adam optimizer, with an initial learning rate of 0.0002 and a weight decay of 0.01. The learning rate follows a one-cycle policy, with a max value of 0.0002 and momentum of [0.95, 0.85]. Training is conducted on a single Nvidia A40 GPU with a batch size of 1 for 24 epochs. A batch size of 1 is also used during inference.
TABLE II: Ablation experiment
Method mIoU Results
Image Point Cloud Average
Baseline 81.32 90.00 85.66
+Mimic Loss 82.90 89.84 86.37 (+0.71)
+Interpolation Attention Feature 84.75 90.48 87.62 (+1.96)
+Depth Map 85.23 91.50 88.37 (+2.71)
+Light Field Image 84.97 92.38 88.68 (+3.02)
IV-B Ablation
TABLE II presents our ablation study results. The baseline extends Mseg3D by incorporating image branch annotation. Our alignment loss enhances image branch performance but slightly degrades the point cloud branch, likely due to partial misalignment caused by occlusion. Despite this, the overall mIoU improves by 0.71. Interpolating the point cloud feature map on the image plane significantly boosts image segmentation, indirectly benefiting the point cloud branch, leading to a 1.96 mIoU increase. Adding the depth difference perception module further enhances occluded object segmentation, raising mIoU by 2.71. Finally, integrating multi-view light field images improves the mIoU by another 2.71. These results validate the effectiveness of the point-pixel interpolation, depth difference perception, and light field integration.
V CONCLUSION
In this work, we introduce the first multimodal dataset for real-world traffic scenes that integrates light field and LiDAR modalities. It includes 623×9 annotated light field images from a 3x3 camera array and corresponding LiDAR point cloud data across 623 frames. Unlike existing datasets, our camera array provides a broader perspective disparity, improving semantic segmentation accuracy. A key feature is the per-pixel annotation across all viewpoints, enabling comprehensive light field segmentation. We benchmark leading single-image, light-field and point-cloud segmentation techniques, showcasing the effectiveness of our dataset. Additionally, we introduce Mipfseg and Mlpfseg, the first multimodal fusion segmentation network for both point clouds and light field images. The network uses point-pixel interpolation fusion to create dense features from sparse point cloud data, with enhanced fusion via alignment loss and cross-attention mechanisms. The depth difference perception module strengthens the learning of occluded object parts, improving recognition accuracy. By incorporating light field images, our approach improves segmentation accuracy for small and occluded objects, achieving state-of-the-art performance for both image and point cloud modalities. In future work, we aim to integrate light field depth estimation for joint training, further refining the depth perception module and developing a more unified network.